### v1

# ArtWhisperer: A Dataset for Characterizing Human-AI Interactions in Artistic Creations

Abstract
As generative AI becomes more prevalent, it is important to study how human users interact with such models. In this work, we investigate how people use text-to-image models to generate desired target images. To study this interaction, we created ArtWhisperer, an online game where users are given a target image and are tasked with iteratively finding a prompt that creates a similar-looking image as the target. Through this game, we recorded over 50,000 human-AI interactions; each interaction corresponds to one text prompt created by a user and the corresponding generated image. The majority of these are repeated interactions where a user iterates to find the best prompt for their target image, making this a unique sequential dataset for studying human-AI collaborations. In an initial analysis of this dataset, we identify several characteristics of prompt interactions and user strategies. People submit diverse prompts and are able to discover a variety of text descriptions that generate similar images. Interestingly, prompt diversity does not decrease as users find better prompts. We further propose to a new metric the study the steerability of AI using our dataset. We define steerability as the expected number of interactions required to adequately complete a task. We estimate this value by fitting a Markov chain for each target task and calculating the expected time to reach an adequate score in the Markov chain. We quantify and compare AI steerability across different types of target images and two different models, finding that images of cities and natural world images are more steerable than artistic and fantasy images. These findings provide insights into human-AI interaction behavior, present a concrete method of assessing AI steerability, and demonstrate the general utility of the ArtWhisperer dataset.
1 Introduction
Direct human interaction with AI models has become widespread following the public release of text-to-text models like GPT-4 [17] and PaLM 2 [1] and text-to-image models like Stable Diffusion [24]. These models have seen rapid interest and adoption in diverse industries including engineering, creative writing, art, education, medicine, and law [6, 16, 9, 4, 22, 3, 28]. A large reason for why this rapid adoption has been possible is that people, with no understanding of how the model works or its limitations, are able to interact with and steer the model to complete a wide variety of tasks.
One of the key challenges for developing these models is aligning their output to human-written input text. This is made especially challenging by the broad domain of use cases as well as the diverse prompting styles of different users. Many approaches can be categorized into the broad label of “prompt engineering” where specific strategies for prompting are used to steer a model [18, 13, 37, 30, 31]. Great success has also been found by fine-tuning models with relatively small datasets to better align a model to follow human instructions [19], to condition the model to respond in a specific style [8], or behave differently to specified prompts [36, 7].
In this work, we take special interest in the fact that human interaction with these models often is an iterative process and develop a dataset to help investigate how people interact with these models with a goal of improving human-AI interaction. To study this interaction, we created an interactive game where players try to find an optimal prompt for a given task (see Figure 1). In particular, we focus on text-to-image models and ask the player to generate a similar image (AI Image) to a given target image. The player is allowed to iterate on their prompt, using the previously generated image(s) as feedback to help them adjust their prompt. A score is also provided as feedback to help the user calibrate how “close” they are to a similar image.
Using this setup, we collected interaction data on 51,026 interactions from 2,250 players across 191 unique target images. The target images were selected from a diverse set of AI-generated and natural images. We also collected a separate dataset of 4,572 interactions, 140 users, and 51 unique target images in a more controlled setting with two different diffusion models to assess the robustness of our findings.
Based on this data, we find several interesting patterns in how people interact with AI models. Players are able to find many different ways to achieve high scores; in particular, they find a diverse set of prompts that all result in a player-generated image similar to the target image. We also find that players tend to make small, iterative updates to their prompts as they steer the AI, with each update improving their image with a moderate success rate ($40-60\%$ for most target images). Based on these findings, we define and evaluate a metric model steerability based on the stopping time of an empirical Markov model. We use this metric to assess steerability across different groups of images and across two AI models.
Ethical considerations
One of the main goals of this work is to help improve the quality of human-AI interaction. our dataset and findings provide quantitative insights on how people interact with generative AI and can potentially be used to design AI that are easier for people to use. It does not address the broader concern that bad actors may abuse generative AI models.
Our contributions
We release a public dataset on human interaction with an AI model. To our knowledge, this is the first such dataset showing repeated interactions of people with a text-to-image model to accomplish specified tasks. We also provide an initial analysis of this data and propose a simple-to-calculate metric for assessing model steerability. Our dataset and associated code is made available at https://github.com/kailas-v/ArtWhisperer.
Figure 1: Interface of ArtWhisperer, our human-AI interaction game. Prompts entered on right. Target image and player-generated image on left. In the lower right is a section where previous prompts and scores are displayed.
Related Works
Datasets for human interaction with text-to-text or text-to-image models typically focus on single interactions (a single user prompt and the resulting model output) and generally do not provide users with a specific task. Public text-to-image interaction datasets typically contain the generated AI images or prompts [26], paired data with both images and prompts [29], and data also paired with some form of human preference rating [20, 34, 35, 11]. These datasets generally rely on scraping online repositories like Lexica [12] or Discord servers focused on AI Art generation. Though some of these datasets include metadata like timestamp and a user ID allowing reconstruction of prompt iteration, there is no guarantee the user has the same target image in mind or whether the user has changed their desired output over time. Public text-to-text datasets are much more limited as the best performing models are generally accessible only through APIs with no public repositories that aggregate user interactions. While some researchers have investigated how human-AI interaction for text-to-text can be improved through various tools [32, 33], the amount data collected is limited and not publicly available. There are also repositories containing prompt strategies for various tasks [2], but no human interaction component.
We seek to rectify two of the shortcomings of the existing datasets–namely, that they do not contain extended interactions as the user attempts to steer the AI, and they do not have a predefined goal. In our work, we create a controlled environment where we allow extended interactions and have a known goal for human users. As shown by our initial analysis, our dataset may enable deeper understanding of user prompting strategies and assessing model steerability.
2 Interaction Game
In the game, players are shown a target image. A few example target images are provided in Figures 24. Players are also provided a limited interface to a text-to-image model; in particular, they are given access to the “positive prompt” and “negative prompt” inputs to the Stable Diffusion (SD) v2.1 model [25]. Here, the positive prompt should contain what the user would like to see in the generated image, while the negative prompt should contain what the user would like omitted from the generated image. Upon inputting a prompt (which often only includes a positive prompt and no negative prompt), the player is shown the image generated by the AI model, along with a score based on similarity between their generated image and the target image. This interface is shown in Figure 1.
Note that players only have access to the model through the positive and negative prompts. In particular, all other parameters are fixed, including the random seed. We opted to fix the random seed for each target image so that all players will generate the same image given the same prompt.
2.1 How Target Images are Selected
We randomly sample target images from two sources. The first is a collection of Wikipedia pages, and the second is a dataset of prompts AI artists have used with SD [26]. In addition to sampling target images, we need to ensure the task is feasible to users. As we do not allow users to adjust the seed or other parameters of the model, we need to ensure the selected model parameters can generate reasonably similar images to the target image. We find that selecting an appropriate random seed is sufficient, and fix all other model parameters (see Section 2.3 for details). Below, we describe how we randomly sample target images and select the random seed.
Wikipedia images
A collection of 35 Wikipedia pages on various topics including art, nature, cities, and various people. A full list of pages sampled from is provided in the Appendix. From these pages, we scraped 670 figures licensed under the Creative Commons license. These figures were then filtered by which had captions, as well as which images were JPG or PNG images (i.e., not animated, and not PDF files), resulting in 557 images.
For each of the 557 images, we first resize and crop the image to size $512\times 512$. The Wikipedia caption is used as the ground truth “prompt”. Let the image-caption pair be denoted as $(t_{i},p_{i}^{*})$. We sample the model on 50 random seeds, with $p_{i}^{*}$ as the prompt input. This generates a set of 50 images: $S_{i}=\{(x_{i},s_{i}):i=1,\dots,50\}$ for generated image $x_{i}$ and seed $s_{i}$. Let $C(x)$ denote the CLIP image embedding [23] of an image $x$. Then we select the seed as $s_{i^{*}}$, where
$i^{*}:=\min_{i=1,\dots,50}\frac{||C(x_{i})-C(t_{i})||_{2}}{||C(x_{i})||_{2}\cdot||C(t_{i})||_{2}}.$
AI-Generated Images
A collection of 2,000 AI-art prompts are randomly sampled from the Stable Diffusion Prompts dataset [26]. For each prompt, $p_{i}^{*}$, we generate two sets of images using 60 unique random seeds: the first set, $S_{i,1}=\{(x_{i},s_{i}):i=1,\dots,10\}$ and $S_{i,2}=\{(x_{i},s_{i}):i=1,\dots,50\}$. We select the target image, $t_{i_{1}^{*}}$, from $S_{i,1}$:
$i_{1}^{*}:=\min_{i=1,\dots,10}\text{median}\left(\left\{\frac{||C(x_{i,1})-C(x_{j,2})||_{2}}{||C(x_{i,1})||_{2}\cdot||C(x_{j,2})||}:j=1,\dots,50\right\}\right)$
We select the random seed, $s_{i_{2}^{*}}$, using $t_{i_{1}^{*}}$ and $S_{i,2}$, with
$i_{2}^{*}:=\min_{i=1,\dots,50}\frac{||C(x_{i,2})-C(t_{i_{1}^{*}})||_{2}}{||C(x_{i,2})||_{2}\cdot||C(t_{i_{1}^{*}})||_{2}}.$
The intuition here is that $t_{i_{1}^{*}}$ is more representative of the types of images we may expect given the fixed prompt, $p_{i}^{*}$.
2.2 Scoring Function
To provide feedback to players, we created a scoring function to assess the similarity of a player’s generated image and the target image. We define the scoring function as
$score(x_{i},t_{i})=\max(0,\min(100,\alpha\cdot\frac{\langle CLIP(x_{i}),CLIP(t_{i})\rangle}{||CLIP(x_{i})||_{2}\cdot||CLIP(t_{i})||_{2}}+\beta)),$
for generated image $x_{i}$, target image $t_{i}$, and constants $\alpha,\beta$. Note the range of $score(x_{i},t_{i})$ is integers in the interval $[0,100]$. Details on how $\alpha,\beta$ are selected parameters are provided in the Appendix.
While this scoring function is often reasonable, it does not always align with the opinions of a human user. To assess how well $score(x_{i},t_{i})$ follows a user’s preferences, we acquire ratings from a subset of users (see ArtWhisperer-Validation in Section 2.4). We find $\text{score}(x_{i},t_{i})$ has a Pearson correlation coefficient of 0.579 indicating reasonable agreement. Further assessment is performed in Section 4.3.
2.3 Additional Technical Details
For the generative model, we use SD v2.1 [25]. We use the DPM Multi-step Scheduler [14], and run the model for 20 iterations. AI-generated target images use the same parameters, but we run the model for 50 iterations. The primary reason for this difference was to limit latency for players; running for 20 iterations resulted in a latency of 1-3 seconds depending on the player’s internet connection. All images are generated at size 512 $\times$ 512.
Figure 2: Example user trajectories. In each row, the first 4 images show the prompt progression for a given user. Target image shown in green column. Plot shows the average score trajectory across all users for this target image (blue) and the specific user’s full score trajectory (red). Orange circles indicate the displayed images.
2.4 Dataset Overview
We collected two datasets as part of this work, which we term ArtWhisperer and ArtWhisperer-Validation. We use ArtWhisperer for most analysis and results; for some of the results in Sections 2.2, 4.2,and 4.3, we also use ArtWhisperer-Validation.
Data was collected from March-May 2023. IRB approval was obtained.
ArtWhisperer: A public version of our game was released online at https://artwhisperer.io/ and we collected data from consenting users playing the game. Three new images were released each day, and users can play during as many days as they would like. Users were anonymous and we only collected data related to the prompts submitted to ensure privacy of users. While we expect some users played the game across multiple days, we did not track them and so do not know the true number of unique users across days. A summary of the ArtWhisperer dataset is provided in Table 1. In total, we have 2,250 players corresponding to 51,026 interactions across 191 target images. We further break down the data across several common categories the images relate to. In Figure 3, we show a density plot of how many queries are submitted by players across different target images.
ArtWhisperer-Validation: A second version of the game (with a near identical interface) was released to paid crowd workers on Prolific [21]. The crowd workers were compensated at a rate of \$$\$12.00$ per hour for roughly 20 minutes of their time. Workers played the game across 5 randomly selected target images from a pre-selected subset of 51 target images chosen to have diverse content. Workers were also asked to rate each of their images on a scale of 1-10 (i.e., self-scoring their generated images). In total, we collected data on 4,572 interactions, corresponding to 140 users and 51 unique target images across two different diffusion models, SD v2.1 and SD v1.5. Additional details on the ArtWhisperer-Validation dataset are provided in the Appendix.
Table 1: ArtWhisperer Dataset Overview. Each row contains summary data for a different subset of the dataset. Subsets may overlap. Similar information for ArtWhisperer-Validation is in the Appendix.
# Players # Target Images # Interactions Average
# Prompts
Average Score Median Duration Category
2250 191 51026 9.29 58.93 18 s Total
377 25 3884 8.65 56.70 19 s Contains famous person?
353 32 3785 8.26 61.64 21 s Contains famous landmark?
2005 140 40290 9.24 59.83 18 s Contains man-made content?
1177 58 18255 10.93 57.21 17 s Contains people?
344 77 6972 8.81 62.01 20 s Is real image?
2140 103 43524 9.42 58.37 17 s Is AI image?
1483 82 24913 9.14 59.45 17 s Is art?
623 29 7297 9.14 53.77 18 s Contains nature?
160 14 1355 7.28 65.74 19 s Contains city?
1239 39 15872 9.91 56.74 16 s Is fantasy?
618 19 8359 10.51 57.88 17 s Is sci-fi or space?
Figure 3: Left, Distribution of # of queries per target image. The average number of queries per image is 9.18. Right, Distribution of the # of words submitted in a query. The average number of words submitted in a positive and negative prompt are 20.02 and 2.32 respectively.
3 Prompt Diversity
Across images, regardless of image content, people submit a diverse set of prompts. Notably, people are able to achieve high scores with a diverse set of prompts. While not necessarily surprising that there are multiple ways to achieve a good score (this indicates that the score in our game has multiple local maxima for any given target image), it is interesting that the high-scoring prompts submitted by users remains diverse. Some examples are shown in Figure 4.
We quantify prompt diversity by looking at the distribution of prompts in the text embedding space. In particular, we use the CLIP text embedding [23], and measure the distance in the embedding space between a given prompt and the average submitted prompt for the corresponding target image. Though the results presented here are based on the CLIP embedding, we do find similar results for other standard text embeddings (see Supplement).
3.1 Diverse prompts used for high scores
In the left of Figure 5, we plot the distribution of distances between the first prompt (in blue) and the last prompt (in orange) to the average prompt for the corresponding target image. Despite the average score improving from 51.9 to 70.3 (out of 100) indicating a significant improvement in score, prompt diversity does not significantly diminish. That is, users do not converge to similar prompts to achieve high scores. Similar analysis of the image embedding space suggests image diversity decreases (see Appendix for details). This makes sense given that the score is linearly related to distance in the CLIP image embedding space. This analysis suggests that, over the course of interaction with the AI model, the distribution of user prompts undergoes a transformation (to be closer to the target image), but does not change in diversity.
3.2 People submit similar prompts throughout their interaction
[…]
While this is not a surprising result (that users often do not make significant changes to their prompt), but it is an important result to understand how typical users interact with AI models. Moreover, it suggests that user initialization (i.e., the first prompt they submit) is critical.
3.3 People have similar prompt styles across images
For each user and each target image, we calculate the average prompt embeddings and the average prompt embedding submitted by the given user. We take the difference of these two embeddings as a user specific average prompt, that represents the style of the user. We then compute the standard deviation these user specific embeddings across target images, as a measure of how much a user varies in prompting style across images.
In the right of Figure 5, we plot the distribution of the standard deviations for users (blue) and for permuted users (orange), where the permuted users are generated by sampling user specific average prompts and assigning them to the simulated players, allowing us to test whether real players have more similarity than random across target images.
[…]
Figure 4: Diverse, high-scoring prompt submissions from players. 4 examples of high-scoring prompts from different users. Target image shown in green column. Plots show image (left) and text (right) embeddings of the displayed images (target in orange; user submissions in green; first submission by corresponding user in red) using the Umap [15] projection.
Figure 5: Left, Diversity of prompts does not diminish as users improve their score with better quality prompts. Center, Individual users have similar prompts for a given image. Right, Individual users have different styles of prompting.
4 Model Steerability
Model steerability refers to the ability of a user to steer a model towards a desired outcome. There is no current consensus on how to measure AI steerability. A common approach is to simply measure performance of a model on standardized dataset evaluations [10, 17]. While this can enable comparisons between tasks and models, this approach does not allow for the feedback loop present when humans interact with a model. Steerability can also be measured qualitatively based on user assessment of their experience interacting with the AI [5]. We create a simple yet informative measure of model steerability. We then analyze this measure across different subgroups of images and across two different Stable Diffusion models–SDv2.1 and the older SDv1.5 [24].
4.1 Measuring steerability
As discussed in Section 3.2, users typically engage with the model through clusters of similar prompts. They typically start with an initial base prompt and proceed to make multiple incremental modifications to it. We use this observation as a basis for creating a steerability metric. We define a Markov chain between scores. Each node is a score with edges connecting to the subsequent score. To make this tractable for empirical analysis, we bin scores into five groups: $[0,20],[21,40],[41,60],[61,80],[81,100]$. We use the expected time taken to reach the last score bin, $[81,100]$, as our steerability score (i.e., the stopping time to reach an adequate score).
For each target image, we calculate the empirical transition probability matrix between binned scores using all the players’ data for that image. We then calculate the steerability score for the given target image by running a Monte Carlo simulation to estimate stopping time, as defined above. To assess steerability across a group of images, we average steerability score across all images in the group.
4.2 Analysis
In Figure 6, we plot the steerability score across image groups. Error bars show the standard error. For examples of steerability scores for individual images, see the Supplement. We find that images containing famous people or landmarks, real images (not AI generated), contain cities, or contain nature are the most steerable. AI-generated images, fantasy images, and images of human art are the least steerable. There are a few possible explanations. The model we are assessing here, SDv2.1, as well as its text encoder OpenCLIP, are trained on subsets of LAION5B [27]. The contents of LAION5B are predominantly real world images, indicating why these images may be more steerable (i.e., text describing these types of images may have a better encoding). Moreover, the prompts for AI-generated images and fantasy images generally include specific internet artists and/or art styles which may not be known to most users making achieving the desired target image more difficult. Another potential reason is the distribution of images chosen for each category. Clearly, there are “easier” and “more difficult” images in each category; part of the reason for smaller stopping time may be the sample of images chosen rather than the actual image category.
We also compare steerability across the two models: SDv2.1 and SDv1.5. Across most image categories, we observe a similar steerability. Images of nature, sci-fi or space, and real images have the largest differences in steerability between the two models; SDv2.1 is more steerable in all three cases. This suggests that SDv2.1 may be more steerable for natural images as well as sci-fi images, and is similarly steerable for other kinds of images including AI-generated artwork. One explanation may be that most of our users were not aware of certain prompting strategies that help models generate more aesthetic images or certain art styles; it is possible that for experienced users, AI art images may be more steerable, and differences between models may be magnified if, for example, a user is experienced working with one particular model. More discussion is provided in the Appendix.
4.3 Justification for automated score
One limitation in our steerability metric comes from the method of scoring user-submitted prompts. Ideally, we would like to assess steerability based on a user’s personal preferences. As mentioned in Section 2.2, the scores and human ratings have a positive correlation. Here, we use the human ratings instead of our score function to assess steerability. We compute the steerability score across both models and across image groups. Generally, the steerability scores change little. In all but two cases (SDv2.1 on sci-fi and space images; SDv1.5 on nature images), the human rating-based steerability score remains within a 95% confidence interval of the score-based steerability score. While our score function may not perfectly capture human preferences, the steerability score we generate appears to be robust to these issues. Further discussion is included in the Appendix.
Figure 6: Steerability across image groups (smaller indicates increased steerability). Bars show average expected stopping time across images in the image group; error bars show standard error of the mean.
5 Discussion
As demonstrated in our initial analysis, the ArtWhisperer dataset can provide insights into user prompting strategies and enables us to assess model steerability for individual tasks and groups of tasks. What makes our dataset particularly useful is the controlled interactive environment, where users work toward a fixed goal, that we capture data in. We expect the ArtWhisperer dataset will be useful beyond the initial analyses presented here. For example, some future work may involve fine-tuning models to better accommodate users based on their prompt trajectories or guide users towards better prompting strategies. There is also deeper analysis possible for understanding user prompting strategies.
Appendix A Appendix
A.1 Information on Wikipedia pages scraped
Table 2 presents a list of the Wikipedia pages used to select real-world target images (see Section 2.1). We extracted images from each listed Wikipedia page. We then uniformly sample a category and subsequently sample an image from a page in that category. This ensures a diverse set of images, which is important given that some of the Wikipedia pages contain many more images than others (e.g., Paris has 10 times more usable images than Social_documentary_photography).
Table 2: Wikipedia images used.
Category Wikipedia Pages
Art Art, Fine_art, Fine-art_photography, History_of_art, Painting
Astro Astrophotography
Buildings Architectural_photography, Architecture, Real_estate
City New_York_City, Paris, San_Francisco, Seoul
Fashion Fashion_design, Fashion_photography, Model_(person)
General Aerial_photography, Culture, Documentary_photography, Photography, Social_documentary_photography
Landscape Landscape_photography
Nature Nature_photography
Plants Flower
Portrait Mug_shot, Portrait_photography, Selfie
US Americans, President_of_the_United_States, United_States
Wildlife Aquatic_ecosystem, Macro_photography, Marine_habitats, Wildlife_observation, Wildlife_photography
A.2 Scoring function details
In Section 2.2, we defined the scoring function, $score(x_{i},t_{i})$ as
$score(x_{i},t_{i})=\max(0,\min(100,\alpha\cdot\frac{\langle CLIP(x_{i}),CLIP(t_{i})\rangle}{||CLIP(x_{i})||_{2}\cdot||CLIP(t_{i})||_{2}}+\beta)).$
$\alpha$ and $\beta$ are constants used to scale the embedding distance prior to clipping the score. To select $\alpha$ and $\beta$, we used a small dataset of interactions collected by the authors prior to the main dataset collection (this data is not included in the released dataset). This dataset contains groups of images paired with scores in the range $[0,1]$. For each target image in this small dataset (5 in total), we add the following images to the dataset:
  1. 1.
    AI-generated images that use the target prompt but with a different seed. These images are assigned a score of $1$.
  2. 2.
    The images corresponding to the human-generated prompts. These images are assigned a score of $0.5$.
  3. 3.
    AI-generated images that use a different prompt than the target prompt. These images are assigned a score of 00.
The intuition here is that with the AI-generated images, we can assume using the same target prompt with a different seed should generate a similar image hence the highest score possible ($1$). Using a different prompt (from our prompt dataset [26]), however, should result in an entirely different image hence the lowest score possible (00). Images generated by people are assumed to be somewhere in between, hence the score of $0.5$.
We then fit a linear regression model to this dataset (to predict score given the CLIP image embedding), using balanced sampling across the image groups. This linear model has parameters
$\alpha^{\prime}$ $=-1.503$
$\beta^{\prime}$ $=1.791$
Since our score range is $[0,100]$ but cosine similarity has a range of $[-1,1]$, we scale the cosine similarity by $100$, resulting in
$\alpha^{\prime\prime}$ $=-150.3$
$\beta^{\prime\prime}$ $=179.1$
We also add a “score adjustment” term that attempts to normalize image difficulty. In particular, for each target image, we compute the un-clipped score for the target image $t_{i}$ and the image generated using the target prompt with, $x_{i}$:
$unclipped\_score(x_{i},t_{i})=\alpha^{\prime\prime}\cdot\frac{\langle CLIP(x_{i}),CLIP(t_{i})\rangle}{||CLIP(x_{i})||_{2}\cdot||CLIP(t_{i})||_{2}}+\beta^{\prime\prime}.$
This score assess the score a user would receive if they exactly entered the target prompt. We fix this value as $100$ (i.e., a perfect score prior to clipping), and set the score adjustment parameter, $c_{i}$, to appropriately normalize this score. In particular,
$c_{i}=\frac{100}{unclipped\_score(x_{i},t_{i})},$
and then we obtain target specific parameters,
$\alpha_{i}$ $=-150.3\cdot c_{i}$
$\beta_{i}$ $=179.1\cdot c_{i}$
We use the target specific parameters, $\alpha_{i}$, $\beta_{i}$, for our scoring function parameters, $\alpha$, $\beta$, (so each target image may have slightly different parameters).
A.3 Additional information on running the game
Game instructions
Game instructions are provided in Figure 7. Here we show the main instructions provided on how to play (top), as well as the tool-tips given for positive prompts (lower left) and negative prompts (lower right).
Figure 7: Game instructions
Crowd workers
Crowd workers are adults from the US. They were paid at a rate of \$$\$12.00$ per hour for roughly 20 minutes of time. Additionally, they were provided bonus payment of between \$$\$0.10-\$0.50$ per image they received a perfect score on (depending on the image difficulty). In total, we paid about $600$ for recruiting the crowd workers.
A.4 Additional example images
We provide additional examples of image trajectories and diverse images in Figures 8 and 9.
Figure 8: More examples of user trajectories, as in Figure 2.
Figure 9: More examples of diverse high-scoring prompts, as in Figure 4.
A.5 Algorithm for steerability
We describe the algorithm for assessing steerability in more detail in Algorithm 1. Here, we define three procedures. EstimateStoppingTime estimates the steerability of a target image. We define a set of score bins; we chose 5 equally sized bins so that there was sufficient data to cover each bin. We also use a regularizer, $\epsilon$. What $\epsilon$ essentially does is encode a prior that from any given score, the transition to a new score is uniformly random. Then for each target image, we find the empirical score transition probabilities in EstimateMarkovChain and use Monte Carlo simulation to estimate the stopping time in RunMonteCarloEstimation.
Input: Set of Images
Output: Steerability estimate
Procedure EstimateStoppingTime(targetImages)
       Initialize array $steerability$ to track stopping times;
       Initialize $scoreBins\leftarrow[[0,20],[21,40],[41,60],[61,80],[81,100]]$;
       Initialize $\epsilon\leftarrow 1$;
[…]
             ${markov}_{i}$leftarrow EstimateMarkovChain(${image}_{i}$, $scoreBins$, $\epsilon$);
             $steerability[i]$leftarrow RunMonteCarloEstimation(${markov}_{i}$);
       end for
      return $\mathbf{E}[steerability]$
Procedure EstimateMarkovChain(targetImage, bins, $\epsilon$)
       Initialize bin pair count as $counts[(\text{bin}_{i},\text{bin}_{j})]\leftarrow\epsilon$, using transitions from a dummy node to model the first prompt submitted;
[…]
                   Convert $\text{score}_{i,j}$ to bin number, $\text{bin}_{i,j}$;
                   Increment $counts[(\text{bin}_{i-1,j},\text{bin}_{i,j})]$ by $1$;
             end for
       end for
      Normalize empirical transition counts to find empirical node transition probabilities;
       Define $markov_{\text{targetImage}}$ using the node transition probabilities;
       return $markov_{\text{targetImage}}$
Procedure RunMonteCarloEstimation(markovChain)
       Run Monte Carlo simulation to estimate time to reach the last bin for markovChain, starting from the dummy (initial) node;
       return Estimated stopping time;
Algorithm 1 Algorithm for estimating image steerability.
A.6 Steerability across models
In Figure 10, we plot the steerability across SDv2.1 and SDv1.5. As described in Section 4.2, images of nature, sci-fi or space, and real images have the largest differences in steerability between the two models. “Nature” is the only image group with a steerability difference greater than the standard deviation of the mean. Other image groups seem to have similar performance across both SDv2.1 and SDv1.5. This suggests that SDv2.1 only makes minor improvement over SDv1.5 across most image categories.
Figure 10: Steerability across image groups, comparing between models using a subset of the images. ”Model 1“ is Stable Diffusion v2.1 and ”Model 2“ is Stable Diffusion v1.5. Note only a subset of the images are used to compare the two models, hence the differences between ”Model 1“ and the results in Figure 6.
A.7 Steerability scores for individual images
In Figure 11, we provide some example images along with their steerability scores. Note that more simple images with well-defined content that likely has high presence in the model’s training data (e.g., the first two rows–a fly on a leaf; a drawing of Barack Obama, well-known public figure) have smaller steerability values, indicating they are easier to steer. However, more complex content that is also more ambiguous for users (e.g., the last three rows), have larger steerability indicating greater difficulty in steering.
Figure 11: Steerability of individual target images. Each row is a different target image. The first four columns are examples of user submitted prompts with the corresponding AI-generated images. Right column (highlighted in green) shows the target image along with the steerability of that image. A smaller steerability value corresponds to the image being easier to steer.
A.8 Additional discussion around human ratings
Here we provide plots for the analysis using human ratings. In Figure 12, we show a scatter plot of scores and human ratings. We also plot a best fit line which has a correlation of 0.597, indicating that our score function produces values that are indeed similar to the human ratings.
We also provide a plot comparing the steerability value calculated using our score function and calculated using the human ratings in Figure 13. Error bars indicate the standard error. Images depicting sci-fi or space have the greatest difference (humans seem to be harsher judges of their generated images’ similarity to the target. However, for most image groups, the two steerability scores are quite close and generally exceed a 95% confidence interval.
Figure 12: Scatter plot of scores and human ratings. Best fit line (linear regression) is shown in black.
Figure 13: Comparing steerability across image groups when computing steerability using the score function and when using the human rating function.